Donny borrows $20 from his brother Mike at a daily interest rate of 5%. How much will he owe his brother after 14 days?
Answer by checkley71(8403)

A=P(1+R)^T
A=20(1+.05)^14
A=20(1.05)^14
A=20*1.98
A=$39.60 AMOUNT OWED AFTER 14 DAYS.